# Chiral phononic and electronic edge modes of EuPtSi

### v1

@@ III.1 Bulk properties
All atoms in EuPtSi are located in the $4a$ Wyckoff position ($x$,$x$,$x$), thus the unit cell contains four chemical formulas. After optimization the crystal structure, we found $a=6.429$, which is in excellent agreement with the experimental value of $a=6.436$ [55]. Additionally, the “free” parameters $x$ describing atomic positions were determined to be $0.3665$, $0.0829$, and $0.6634$ for Eu, Pt, and Si, respectively, while the experimentally reported values are $0.3700$, $0.0863$, and $0.6611$ [55].
@@
Phonon spectra.— The phonon dispersion curves are presented in Fig. 2. Due to the absence of imaginary soft modes, EuPtSi is stable in the P213 structure in dynamical sense. In the phonon spectra, we can recognize a few groups of bands: (i) a low-frequency range (below $5$ THz), containing, e.g., acoustic branches; (ii) a middle-frequency range (around $6$ THz); and (iii) a high frequency range (around $12$ THz). The acoustic branches around the $\Gamma$ point possess nearly ideal linear dispersion and are well hybridized with optical branches above $2$ THz. Optical branches in the middle and high-frequency ranges are relatively narrow.
@@
The formation of such groups of bands is also reflected in the phonon density of states (phDOS), presented in Fig. 2(b). Due to their atomic masses, the vibrations associated with the heavies atoms (Eu and Pt, with comparable atomic masses) occur mostly in low-frequency range. In practice, the small contribution of Si vibrations to the phDOS in the low-frequency range is realized only within the acoustic branches. Moreover, the middle and high-frequency phonon modes are associated with vibrations of Si atoms. Above $5$ THz, there is only small contribution of Pt atoms to the phDOS (at the top of the middle-frequency group and the bottom of the high-frequency group). This is associated with the vibration of Pt and Si atoms, which are well coupled within the same Pt-Si chains. A similar situation was observed in BaPtGe, also with P213 symmetry, where high-frequency branches (with Pt and Ge contributions) are separated from another modes (mixed by vibrations of all atoms) [54].
@@
The crystal symmetries gives rise to threefold representations at the $\Gamma$ point and fourfold representations at the R point [see inset in Fig. 2(a)]. As consequence, in the phononic band structure, we can recognize a spin-1 Weyl point (at $\Gamma$) and an effective charge-2 Dirac point (at R), which are special cases of double Weyl points [51]. Both topological points possess opposite non-zero Chern number (±$\pm 2$), and their total Chern numbers cancel exactly within the Brillouin zone. The spin-1 Weyl point is formed by the crossing of three branches, where two of them form Dirac cones, while the third forms a nearly-flat band. The crossing of all these bands at the same frequency gives rise to the triple degeneracy at the $\Gamma$ point. In contrast, the charge-2 Dirac point is formed directly by the fourfold degenerate states at R. In this case, a doubly-degenerate Dirac cone is realized around the R point. Both points, $\Gamma$ and R, exist due to the crystal symmetry and time-reversal symmetry. In practice, the spin-1 Weyl and charge-2 Dirac points can be found in the phonon spectra over a wide range of frequencies.
@@
Electronic and and Fermi surface.— The electronic band structure for Eu $f$ electrons treated as core and valence states are presented in Fig. 3(a) and 3(b), respectively. As we can see, the localized Eu $f$-states are located far below the Fermi level (around $-1.75$ eV). When $f$-states are treated as valence states, we observe additional band splitting due to the exchange field acting on the other states. However, the $f$-states do not strongly affect the electronic band structure [cf. blue lines in Fig. 3(a) and 3(b)]. Moreover, the electronic band structure of EuPtSi with $f$ electrons treated as a core states is similar to that reported earlier for SrPtSi (without $f$-states) [65].
@@
In the absence of SOC, the electronic band structure features a spin-1 Weyl point and an effective charge-2 Dirac point at the $\Gamma$ and R points, respectively [e.g., orange lines in the inset of Fig. 3(a)]. For instance, the spin-1 Weyl point is visible around $-2.1$ eV, while the charge-2 Dirac point is around $-2.75$ eV. Introducing SOC leads to the decoupling of spin-1 state from a spin-degenerate orbital triplet into a doubly-degenerate point and fourfold-degenerate states (with a Chern number of $+4$). Similarly, at the R point, SOC leads to the decoupling into a sixfold fermionic point (with a Chern number of $-4$) and a trivial doubly-degenerate point. Such effect is shown in the inset of Fig. 3(a).
@@
We should also note that relatively strong spin–orbit effects [cf. orange and blue lines in Fig. 3(a) and 3(b)] are observed well below the Fermi level (below $-2$ eV). The density of states [Fig. 3(c)] uncover reveals a strong Pt $d$-states contribution in this energy range. Heavy atoms introduce strong SOC, and indeed, Pt has the largest spin–obit contribution within $d$-states ($60$ meV) and $p$-states ($38$ meV). In contrast, for Eu, SOC contributions are $140$ meV, $0.3$ meV, $13$ meV for $p$-, $d$-, and $f$-states, respectively, while spin–orbit affect of Si is negligible ($<0.2$ meV).
@@
The density of states is mainly due to contributions from Eu $d$$f$, Pt $d$- (mostly below $-2$ eV), and Si $p$-states. It is noteworthy that there are no Eu $d$ electrons in initial calculations (Eu has a valence configuration of $5p^{6}6s^{2}$ when $f$-states are treated as core states), but Eu $d$ electrons contribute to the density of states at the Fermi level in the final electronic state.
@@
For simplicity, in our calculations containing the Eu $f$-states, we assumed an antiferromagnetic (AFM) order. Nevertheless, the experimentally observed helimagnetic order should not change the band structure significantly. The experimentally observed saturation magnetic moment is close to the nominal value of $7$ $\mu_{\text{B}}$ [55], as expected for Eu+2 with spin $7/2$ [95, 96]. From the self-consistent calculations, we determined the Eu magnetic moment to be $6.96$ $\mu_{\text{B}}$. The occupied Eu $f$-states are located far below the Fermi level [from $-2.0$ eV to $-1.5$ eV, cf. Fig. 3(a) and 3(b)]. As consequence, the main features of the electronic band structure are the same independently of the Eu $f$ electrons.
@@
Figure 6: The electronic spectral function for the (001) surface along the path presented in Fig. 5(b). The constant energy contour for two cuts shown in (a): cut 1 at $-2$ eV (b), and cut 2 at $0.84$ eV (c). The spectral function along loop 1 and loop 2 (d), around the $\bar{\Gamma}$ and $\bar{\text{M}}$ points, respectively.
@@ III.2 Phononic surface states
Our results regarding the phonon surface states for the (001) surface are presented in Fig. 5. The chiral phononic edge modes should connect the $\Gamma$ and R points of the bulk 3D Brillouin zone projected onto the surface two-dimensional (2D) Brillouin zone, as shown in Fig. 3(b). The bulk $\Gamma$ and R points are projected onto the surface $\bar{\Gamma}$ and $\bar{\text{M}}$ points, respectively. All surface states should be visible as states crossing any path between $\bar{\Gamma}$ and $\bar{\text{M}}$ points, e.g., the $\bar{\text{X}}$¯¯$\bar{\text{Y}}$ path [Fig. 5(b)]. Indeed, the spectral function clearly shows surface states crossing paths close to $\bar{\Gamma}$ and $\bar{\text{M}}$ points, such as the $\bar{\text{X}}$¯¯$\bar{\text{Y}}$’¯¯$\bar{\text{X'}}$’¯¯$\bar{\text{Y'}}$¯¯$\bar{\text{X}}$ path [Fig. 5(a)]. In practice, the chiral edge modes can be found in the whole range of frequencies, in the form of separate states connecting the two (bottom and top) “bands”, for example, the states marked by white arrows in Fig. 5(a).
We also present the path $\bar{\text{X}}$¯¯$\bar{\text{M}}$$\bar{\Gamma}$. As we can see, at $\bar{\Gamma}$ and $\bar{\text{M}}$ we find Dirac-cone-like structures at the frequencies marked by red arrows in Fig. 5(a). Indeed, these points contain states from the highly degenerate bulk points $\Gamma$ and R. In the low- and high-frequency ranges, the Dirac-cone-like crossings are realized in close pairs of frequencies (i.e., $12.08$ THz and $12.00$ THz, $3.96$ THz and $3.86$ THz). Contrary to this, in the intermediate frequencies, both points coincide at the same frequency of $6.10$ THz.
@@ III.3 Electronic surface states
Now, we will discuss the corresponding surface states problem in the context of the electronic spectrum. Contrary to the phonon spectrum, the electronic band structure exhibits additional band splitting due to the strong SOC. As we mention earlier, the SOC leads to decoupling of the spin-1 Weyl and charge-2 Dirac points. The Weyl point decouples into a doubly degenerate point and fourfold degenerats states (with Chern number $+4$), while the Dirac point decouples into a sixfold fermionic point (with Chern number $-4$) and a trivial doubly degenerate point.
Compared to the phonon dispersion curves, the electronic band structure has a much more complex form. Moreover, the SOC-induced decoupling also affects the chiral edge modes. Results of the spectral function obtained for the electronic spectrum of EuPtSi for (001) are presented in Fig. 6. Along the $\bar{\text{X}}$¯¯$\bar{\text{Y}}$’¯¯$\bar{\text{X'}}$’¯¯$\bar{\text{Y'}}$¯¯$\bar{\text{X}}$ path, several surface states are visible [see Fig. 6(a)]. However, not all surface states exhibit chiral features. For example, a chiral edge mode is well visible much below the Fermi level (around $-2$ eV). Moreover, for other energies, the chiral edge mode can be mixed with other edge modes. To illustrate both cases, we show constant energy cross-sections at energies of $-2$ eV and $0.83$ eV [cuts 1 and 2 in Fig. 6(a)]. Results for cut 1 and 2 are presented in Fig. 6(b) and 6(c), respectively. For energy $-2$ eV (cut 1), clearly decoupled chiral edge modes connect the $\bar{\Gamma}$ and $\bar{\text{M}}$ points. Along the Fermi arc, the SOC-induced unconventional spin texture is evident, as indicated by the red arrows in Fig. 6(b). For energy $0.83$ eV (cut 2), the chiral edge mode is less prominent. Surface states are also observed around $\bar{\text{X}}$ and $\bar{\text{Y}}$ [Fig. 6(c)].
@@
Figure 7: Side view (a) and view along the (111) direction (b) of the EuPtSi crystal, together with the calculated charge density distribution for states within energy window from $-2.1$ eV to $-1.7$ eV [marked by red dashed lines in panel Fig. 6(a)].
@@ IV Summary
In the phononic bulk dispersion curves, due to the absence of SOC-like effects, ideal spin-1 Weyl and charge-2 Dirac points are realized at the $\Gamma$ and R points, respectively. This is possible as a consequence of the coexistence of three twofold screw rotations along Cartesian directions and one threefold rotational symmetry axis around the (111) direction. Moreover, the non-zero Chern numbers ±$\pm 2$ at the corresponding points of the 2D surface Brillouin zone (i.e., $\bar{\Gamma}$ and $\bar{\text{R}}$ points) allow for the realization of the chiral edge modes connecting these points. For the constant frequency cross-section, such a chiral edge mode is visible in the form of a Fermi arc, connecting $\bar{\Gamma}$ and $\bar{\text{R}}$ points. Similarly, the chiral edge mode is visible in the phonon spectrum as a mode propagating along one direction in closed loop around the $\bar{\Gamma}$ or $\bar{\text{R}}$ point.
The situation is more complicated in the case of the electronic spectrum. Due to the SOC effect, the ideal spin-1 Weyl and charge-2 Dirac points are split and no longer observed. However, there are still points with non-zero Chern numbers ±$\pm 4$ at $\bar{\Gamma}$ and $\bar{\text{R}}$ points. From this, the edge states are realized in the form of split chiral edge modes. The main features of such modes are the same as in the case of the phononic chiral edge modes. Additionally, an unconventional spin texture exists on the surface states.
@@
Similarly, in the case of the electronic surface state, some states localization can be expected. Indeed, this is well visible in the calculated charge density distribution for states within the energy window from $-2.1$ eV to $-1.7$ eV [marked by red dashed lines in panel Fig. 6(a)], i.e., for the range of energies corresponding to the electronic chiral edge modes. From this, we can see directly that carrier accumulation is realized within Pt-Si chains, only on atoms around the surface. The absence of localization on Eu atoms can be consequence of the relatively large distance between atoms inside the Eu chain. Distances between atoms in Pt-Si chains are around $2.33$, while in Eu chains, they are around $\sim 3.94$. The chiral surface states around the Fermi level do not exhibit a tendency for strong localization around the surface due to the strong mixing with the bulk states. However, the corresponding states still remain along Pt-Si chains.

### v2

@@ III.1 Bulk properties
All atoms in EuPtSi are located in the $4a$ Wyckoff position ($x$,$x$,$x$), thus the unit cell contains four chemical formulas. After optimization the crystal structure, we found $a=6.429$ Å, which is in excellent agreement with the experimental value of $a=6.436$ Å [55]. Additionally, the “free” parameters $x$ describing atomic positions were determined to be $0.3665$, $0.0829$, and $0.6634$ for Eu, Pt, and Si, respectively, while the experimentally reported values are $0.3700$, $0.0863$, and $0.6611$ [55].
@@
Phonon spectra.— The phonon dispersion curves are presented in Fig. 2. Due to the absence of imaginary soft modes, EuPtSi is stable in the P213 structure in dynamical sense. In the phonon spectra, we can recognize a few groups of bands: (i) a low-frequency range (below 55 THz), containing, e.g., acoustic branches; (ii) a middle-frequency range (around 66 THz); and (iii) a high frequency range (around $12$ THz). The acoustic branches around the $\Gamma$ point possess nearly ideal linear dispersion and are well hybridized with optical branches above 22 THz. Optical branches in the middle and high-frequency ranges are relatively narrow.
@@
The formation of such groups of bands is also reflected in the phonon density of states (phDOS), presented in Fig. 2(b). Due to their atomic masses, the vibrations associated with the heavies atoms (Eu and Pt, with comparable atomic masses) occur mostly in low-frequency range. In practice, the small contribution of Si vibrations to the phDOS in the low-frequency range is realized only within the acoustic branches. Moreover, the middle and high-frequency phonon modes are associated with vibrations of Si atoms. Above 55 THz, there is only small contribution of Pt atoms to the phDOS (at the top of the middle-frequency group and the bottom of the high-frequency group). This is associated with the vibration of Pt and Si atoms, which are well coupled within the same Pt-Si chains. A similar situation was observed in BaPtGe, also with P213 symmetry, where high-frequency branches (with Pt and Ge contributions) are separated from another modes (mixed by vibrations of all atoms) [54].
@@
The crystal symmetries gives rise to threefold representations at the $\Gamma$ point and fourfold representations at the R point [see inset in Fig. 2(a)]. As consequence, in the phononic band structure, we can recognize a spin-1 Weyl point (at $\Gamma$) and an effective charge-2 Dirac point (at R), which are special cases of double Weyl points [51]. Both topological points possess opposite non-zero Chern number ($\pm 2$), and their total Chern numbers cancel exactly within the Brillouin zone. The spin-1 Weyl point is formed by the crossing of three branches, where two of them form Dirac cones, while the third forms a nearly-flat band. The crossing of all these bands at the same frequency gives rise to the triple degeneracy at the $\Gamma$ point. In contrast, the charge-2 Dirac point is formed directly by the fourfold degenerate states at R. In this case, a doubly-degenerate Dirac cone is realized around the R point. Both points, $\Gamma$ and R, exist due to the crystal symmetry and time-reversal symmetry. In practice, the spin-1 Weyl and charge-2 Dirac points can be found in the phonon spectra over a wide range of frequencies.
@@
Electronic and and Fermi surface.— The electronic band structure for Eu $f$ electrons treated as core and valence states are presented in Fig. 3(a) and 3(b), respectively. As we can see, the localized Eu $f$-states are located far below the Fermi level (around 1.75-1.75 eV). When $f$-states are treated as valence states, we observe additional band splitting due to the exchange field acting on the other states. However, the $f$-states do not strongly affect the electronic band structure [cf. blue lines in Fig. 3(a) and 3(b)]. Moreover, the electronic band structure of EuPtSi with $f$ electrons treated as a core states is similar to that reported earlier for SrPtSi (without $f$-states) [65].
@@
In the absence of SOC, the electronic band structure features a spin-1 Weyl point and an effective charge-2 Dirac point at the $\Gamma$ and R points, respectively [e.g., orange lines in the inset of Fig. 3(a)]. For instance, the spin-1 Weyl point is visible around 2.1-2.1 eV, while the charge-2 Dirac point is around 2.75-2.75 eV. Introducing SOC leads to the decoupling of spin-1 state from a spin-degenerate orbital triplet into a doubly-degenerate point and fourfold-degenerate states (with a Chern number of +4+4). Similarly, at the R point, SOC leads to the decoupling into a sixfold fermionic point (with a Chern number of 4-4) and a trivial doubly-degenerate point. Such effect is shown in the inset of Fig. 3(a).
@@
We should also note that relatively strong spin–orbit effects [cf. orange and blue lines in Fig. 3(a) and 3(b)] are observed well below the Fermi level (below 2-2 eV). The density of states [Fig. 3(c)] uncover reveals a strong Pt $d$-states contribution in this energy range. Heavy atoms introduce strong SOC, and indeed, Pt has the largest spin–obit contribution within $d$-states ($60$ meV) and $p$-states ($38$ meV). In contrast, for Eu, SOC contributions are $140$ meV, $0.3$ meV, $13$ meV for $p$-, $d$-, and $f$-states, respectively, while spin–orbit affect of Si is negligible (<0.2<0.2 meV).
@@
The density of states is mainly due to contributions from Eu ddff, Pt $d$- (mostly below 2-2 eV), and Si $p$-states. It is noteworthy that there are no Eu $d$ electrons in initial calculations (Eu has a valence configuration of $5p^{6}6s^{2}$ when $f$-states are treated as core states), but Eu $d$ electrons contribute to the density of states at the Fermi level in the final electronic state.
@@
For simplicity, in our calculations containing the Eu $f$-states, we assumed an antiferromagnetic (AFM) order. Nevertheless, the experimentally observed helimagnetic order should not change the band structure significantly. The experimentally observed saturation magnetic moment is close to the nominal value of 77 $\mu_{\text{B}}$ [55], as expected for Eu+2 with spin $7/2$ [95, 96]. From the self-consistent calculations, we determined the Eu magnetic moment to be $6.96$ $\mu_{\text{B}}$. The occupied Eu $f$-states are located far below the Fermi level [from 2.0-2.0 eV to 1.5-1.5 eV, cf. Fig. 3(a) and 3(b)]. As consequence, the main features of the electronic band structure are the same independently of the Eu $f$ electrons.
@@
Figure 6: The electronic spectral function for the (001) surface along the path presented in Fig. 5(b). The constant energy contour for two cuts shown in (a): cut 1 at 2-2 eV (b), and cut 2 at $0.84$ eV (c). The spectral function along loop 1 and loop 2 (d), around the $\bar{\Gamma}$ and $\bar{\text{M}}$ points, respectively.
@@ III.2 Phononic surface states
Our results regarding the phonon surface states for the (001) surface are presented in Fig. 5. The chiral phononic edge modes should connect the $\Gamma$ and R points of the bulk 3D Brillouin zone projected onto the surface two-dimensional (2D) Brillouin zone, as shown in Fig. 3(b). The bulk $\Gamma$ and R points are projected onto the surface $\bar{\Gamma}$ and $\bar{\text{M}}$ points, respectively. All surface states should be visible as states crossing any path between $\bar{\Gamma}$ and $\bar{\text{M}}$ points, e.g., the $\bar{\text{X}}$$\bar{\text{Y}}$ path [Fig. 5(b)]. Indeed, the spectral function clearly shows surface states crossing paths close to $\bar{\Gamma}$ and $\bar{\text{M}}$ points, such as the $\bar{\text{X}}$$\bar{\text{Y}}$$\bar{\text{X'}}$$\bar{\text{Y'}}$$\bar{\text{X}}$ path [Fig. 5(a)]. In practice, the chiral edge modes can be found in the whole range of frequencies, in the form of separate states connecting the two (bottom and top) “bands”, for example, the states marked by white arrows in Fig. 5(a).
We also present the path $\bar{\text{X}}$$\bar{\text{M}}$$\bar{\Gamma}$. As we can see, at $\bar{\Gamma}$ and $\bar{\text{M}}$ we find Dirac-cone-like structures at the frequencies marked by red arrows in Fig. 5(a). Indeed, these points contain states from the highly degenerate bulk points $\Gamma$ and R. In the low- and high-frequency ranges, the Dirac-cone-like crossings are realized in close pairs of frequencies (i.e., $12.08$ THz and $12.00$ THz, $3.96$ THz and $3.86$ THz). Contrary to this, in the intermediate frequencies, both points coincide at the same frequency of $6.10$ THz.
@@ III.3 Electronic surface states
Now, we will discuss the corresponding surface states problem in the context of the electronic spectrum. Contrary to the phonon spectrum, the electronic band structure exhibits additional band splitting due to the strong SOC. As we mention earlier, the SOC leads to decoupling of the spin-1 Weyl and charge-2 Dirac points. The Weyl point decouples into a doubly degenerate point and fourfold degenerats states (with Chern number +4+4), while the Dirac point decouples into a sixfold fermionic point (with Chern number 4-4) and a trivial doubly degenerate point.
Compared to the phonon dispersion curves, the electronic band structure has a much more complex form. Moreover, the SOC-induced decoupling also affects the chiral edge modes. Results of the spectral function obtained for the electronic spectrum of EuPtSi for (001) are presented in Fig. 6. Along the $\bar{\text{X}}$$\bar{\text{Y}}$$\bar{\text{X'}}$$\bar{\text{Y'}}$$\bar{\text{X}}$ path, several surface states are visible [see Fig. 6(a)]. However, not all surface states exhibit chiral features. For example, a chiral edge mode is well visible much below the Fermi level (around 2-2 eV). Moreover, for other energies, the chiral edge mode can be mixed with other edge modes. To illustrate both cases, we show constant energy cross-sections at energies of 2-2 eV and $0.83$ eV [cuts 1 and 2 in Fig. 6(a)]. Results for cut 1 and 2 are presented in Fig. 6(b) and 6(c), respectively. For energy 2-2 eV (cut 1), clearly decoupled chiral edge modes connect the $\bar{\Gamma}$ and $\bar{\text{M}}$ points. Along the Fermi arc, the SOC-induced unconventional spin texture is evident, as indicated by the red arrows in Fig. 6(b). For energy $0.83$ eV (cut 2), the chiral edge mode is less prominent. Surface states are also observed around $\bar{\text{X}}$ and $\bar{\text{Y}}$ [Fig. 6(c)].
@@
Figure 7: Side view (a) and view along the (111) direction (b) of the EuPtSi crystal, together with the calculated charge density distribution for states within energy window from 2.1-2.1 eV to 1.7-1.7 eV [marked by red dashed lines in panel Fig. 6(a)].
@@ IV Summary
In the phononic bulk dispersion curves, due to the absence of SOC-like effects, ideal spin-1 Weyl and charge-2 Dirac points are realized at the $\Gamma$ and R points, respectively. This is possible as a consequence of the coexistence of three twofold screw rotations along Cartesian directions and one threefold rotational symmetry axis around the (111) direction. Moreover, the non-zero Chern numbers $\pm 2$ at the corresponding points of the 2D surface Brillouin zone (i.e., $\bar{\Gamma}$ and $\bar{\text{R}}$ points) allow for the realization of the chiral edge modes connecting these points. For the constant frequency cross-section, such a chiral edge mode is visible in the form of a Fermi arc, connecting $\bar{\Gamma}$ and $\bar{\text{R}}$ points. Similarly, the chiral edge mode is visible in the phonon spectrum as a mode propagating along one direction in closed loop around the $\bar{\Gamma}$ or $\bar{\text{R}}$ point.
The situation is more complicated in the case of the electronic spectrum. Due to the SOC effect, the ideal spin-1 Weyl and charge-2 Dirac points are split and no longer observed. However, there are still points with non-zero Chern numbers $\pm 4$ at $\bar{\Gamma}$ and $\bar{\text{R}}$ points. From this, the edge states are realized in the form of split chiral edge modes. The main features of such modes are the same as in the case of the phononic chiral edge modes. Additionally, an unconventional spin texture exists on the surface states.
@@
Similarly, in the case of the electronic surface state, some states localization can be expected. Indeed, this is well visible in the calculated charge density distribution for states within the energy window from 2.1-2.1 eV to 1.7-1.7 eV [marked by red dashed lines in panel Fig. 6(a)], i.e., for the range of energies corresponding to the electronic chiral edge modes. From this, we can see directly that carrier accumulation is realized within Pt-Si chains, only on atoms around the surface. The absence of localization on Eu atoms can be consequence of the relatively large distance between atoms inside the Eu chain. Distances between atoms in Pt-Si chains are around $2.33$ Å, while in Eu chains, they are around $\sim 3.94$ Å. The chiral surface states around the Fermi level do not exhibit a tendency for strong localization around the surface due to the strong mixing with the bulk states. However, the corresponding states still remain along Pt-Si chains.